### v1

# Go Beyond Earth: Understanding Human Actions and Scenes in Microgravity Environments

Abstract
Despite substantial progress in video understanding, most existing datasets are limited to Earth’s gravitational conditions. However, microgravity alters human motion, interactions, and visual semantics, revealing a critical gap for real-world vision systems. This presents a challenge for domain-robust video understanding in safety-critical space applications. To address this, we introduce MicroG-4M, the first benchmark for spatio-temporal and semantic understanding of human activities in microgravity. Constructed from real-world space missions and cinematic simulations, the dataset includes 4,759 clips covering 50 actions, 1,238 context-rich captions, and over 7,000 question–answer pairs on astronaut activities and scene understanding. MicroG-4M aims to support three core tasks: fine-grained multi-label action recognition, temporal video captioning, and visual question answering, thereby enabling a comprehensive evaluation of both spatial localization and semantic reasoning in microgravity contexts. We establish baselines using state-of-the-art models. All data, annotations, and code are available at https://github.com/LEI-QI-233/HAR-in-Space.
1 Introduction
Figure 1: An illustration of the MicroG-4M dataset, which contains videos from real and simulated microgravity environments (e.g., movies). The dataset supports benchmarks for three tasks: (1) video captioning, (2) video-based question answering, and (3) fine-grained human action recognition under microgravity.
Yuri Gagarin’s historic flight in 1961 marked the beginning of human space exploration. Since then, significant milestones have been achieved, including crewed lunar landings, the continuous operation of the International Space Station (ISS) for over $25$ years, and the participation of more than $650$ individuals in space missions [1]. With numerous planned crewed missions in the near future, the frequency and complexity of human activities in space are expected to increase substantially. In this context, ensuring the safety, enhancing the operational efficiency of space missions, and safeguarding the health and well-being of astronauts are of paramount importance.
With the rapid advancement of artificial intelligence, the integration of robotic systems aboard spacecraft is anticipated in the near future. These systems will assist astronauts in routine and mission-critical tasks. Consequently, there is a growing demand for the development of deep learning-based scene understanding and action recognition methods tailored specifically for the unique challenges posed by the microgravity environment. Human action recognition [2, 3, 4], scenario captioning [5], and Visual Question Answering (VQA) [6] are essential for intelligent human-robot collaboration, particularly in space, where precise perception and understanding of astronauts’ actions and their surrounding context are crucial for ensuring operational safety, efficiency, and autonomous assistance under constrained conditions. This capability is crucial for ensuring mission efficiency, enhancing astronaut safety, and providing autonomous assistance in the confined and complex conditions of space habitats. Numerous human action recognition datasets and video caption datasets have been developed, playing an important role in various research areas within computer vision and in industrial applications [7, 8, 9, 10].
However, most of the existing datasets for video scenario captioning and action recognition are recorded on Earth without microgravity settings, e.g., Kinetics400 [11], AVA [12], FineGym [13], and ActivityNet Captions [14].
Human actions in space differ significantly from those on Earth due to microgravity, where basic movements like standing, walking, eating, and object manipulation follow entirely different physical principles [15]. For instance, ‘standing’ can occur at any angle, ‘walking’ resembles swimming, and objects are often grabbed mid-air rather than picked up or put down conventionally. As a result, existing Earth-based datasets fail to capture these nuances, limiting the effectiveness of Human Action Recognition (HAR) in microgravity environments.
To address this gap, we introduce MicroG-4M, a new video benchmark specifically designed for spatio-temporal and semantic understanding of human activities in microgravity. The name “4M” reflects four core characteristics of the dataset: Multi-domain, Multimodal, Multi-task, and Microgravity. MicroG-4M comprises $4,759$ three-second video clips drawn from public YouTube footage of real space missions and carefully selected, realistic space-themed films to augment scenario diversity and coverage. The clips span scenes inside the ISS, the Tiangong Space Station, crewed spacecraft cabins, and extravehicular activities. The corpus contains more than $390,000$ annotated frames with bounding boxes for each visible individual and $13,000+$ action labels covering $50$ distinct actions. In addition, it includes $1,238$ descriptive clip captions created by human annotators and $7,000+$ open-ended question–answer pairs aimed at assessing factual, causal, and counterfactual understanding of scenes in micro-gravity. Both captions and QA annotations were carefully curated through multiple refinement rounds and data selection stages, ensuring high semantic fidelity, contextual accuracy, and relevance to the microgravity setting. Together, these resources enable multi-label spatio-temporal detection, fine-grained action recognition, caption generation, and VQA within a single dataset. To evaluate existing methods, we build the first comprehensive benchmark for these tasks, dubbed MicroG-Bench. We evaluated representative video encoders (SlowFast [4], X3D [3], and MViT [16]) for human action recognition and leading vision-language models (e.g., InternVideo [17], Gemini 1.5 Pro [18], GPT-4o [19]) for captioning and VQA tasks. Across all evaluated subtasks, state-of-the-art terrestrial models experience significant performance degradation, underscoring the unique challenges posed by micro-gravity environments, such as arbitrary orientations, floating objects, and confined spacecraft interiors. These results highlight the necessity of dedicated benchmarks to facilitate the advancement of more robust and generalizable AI systems tailored for space applications. MicroG-Bench, therefore, provides a unified, fair yardstick that will guide the development of more robust and generalizable perception and language systems for astronaut assistance and autonomous mission operations. Then we present a short summarization of the contribution in this work. We present MicroG-4M, the first video-based dataset for activity recognition and scene understanding specifically designed for the micro-gravity environment. Additionally, we establish the first benchmark to evaluate state-of-the-art human action recognition methods and vision-language models for fine-grained action recognition and video captioning/QA in this unique setting.
2 Related Work
Vision-Based Understanding in Microgravity and Space Environments. Microgravity challenges core assumptions of terrestrial vision systems, necessitating perception models tailored to space habitats. Early work focused on visual localization, where traditional SLAM proved unreliable, prompting robust alternatives using visual-inertial fusion, semantic mapping, and CAD-informed constraints [20, 21, 22, 23], validated on platforms like Astrobee [24]. Real-time vision has also enabled dynamic tasks such as target tracking and scene change detection [25, 26].
As research shifted toward interaction, studies explored astronaut pose recovery [27, 28], gesture recognition [29, 30], EMG-based input [31], and long-term astronaut identification [32]. However, high-level semantic understanding—such as reasoning about astronaut actions and intent—remains limited. Current assistants like CIMON [33] offer basic perception but lack contextual reasoning. To fill this gap, we introduce a unified benchmark for action localization, video captioning, and visual question answering in operational space environments.
Datasets for Action Detection, Video Captioning, and VQA. Progress in video understanding has been largely enabled by the introduction of benchmark datasets across action detection, captioning, and VQA. For action detection, early datasets such as UCF101 [34] and HMDB51 [35] provided trimmed classification tasks, later extended by Kinetics [11] and ActivityNet [36] to large-scale, untrimmed settings with diverse action categories. AVA [12] further introduced spatio-temporal annotations for atomic actions, enabling fine-grained multi-label detection in realistic scenes. Video captioning evolved from MSVD [37] and MSR-VTT [38], which paired short clips with multiple sentences, to dense captioning in untrimmed videos via ActivityNet Captions [39] and procedural datasets like YouCook2 [40]. Multilingual benchmarks such as VATEX [41] expanded the task to cross-lingual settings with high-quality parallel annotations. In VQA, static image datasets like VQA v2.0 [42] and CLEVR [43] laid the foundation for compositional reasoning, while TGIF-QA [44], MovieQA [45], and TVQA [46] introduced temporal and multimodal reasoning in video. Recent efforts such as NExT-QA [47] and CLEVRER [48] focus on causal inference and counterfactual reasoning. However, existing datasets are constrained to terrestrial environments and lack domain-specific complexity inherent to space-based activities. To bridge this gap, we introduce the first benchmark for video captioning and VQA in microgravity, enabling the evaluation of vision-language models in long-horizon, space-relevant scenarios.
Fine-Grained Human Action Recognition. Fine-grained video-based action recognition [49, 50] aims to detect fine-grained, indivisible human actions in both single- and multi-person videos. Unlike standard clip-level recognition, atomic action localization requires frame-level, multi-label classification with spatio-temporal bounding box predictions. To tackle this, CNN-based [2, 3, 4] and transformer-based [51, 2, 16, 52, 17, 53, 49] models have been adapted by adding multi-label heads, bounding box regressors, and region-of-interest modules. Notably, Ryali et al.[51] introduced a hierarchical vision transformer balancing accuracy and efficiency, while Wanget al. [2] proposed a dual-masked autoencoder for improved video pretraining. In this work, we evaluate such HAR methods for the first time in microgravity scenarios.
3 Collection Methodology
We aim to construct a dataset comprising authentic footage recorded aboard the International Space Station and other spacecraft, as well as films that realistically depict microgravity conditions. In the following, we introduce our comprehensive pipeline for collecting, assembling, filtering, and annotating the MicroG-4M dataset, which supports multi-label spatio-temporal action detection, video captioning, and video question answering tasks from Internet-sourced videos.
Raw Video Information Collection. Video sources primarily include genuine microgravity footage from actual spacecraft missions and selected cinematic clips known for their realistic depictions of weightlessness. Authentic spacecraft videos were mainly retrieved from online video platforms, while cinematic clips were manually chosen based on their fidelity to real microgravity conditions. All videos were standardized to a resolution of $480p$ to maintain consistency. The final dataset comprises approximately $5,000$ three-second clips, predominantly composed of authentic spacecraft footage.
Dataset Assembly Pipeline. We established an automated pipeline to preprocess and structure raw videos, enabling consistent and accurate downstream annotation. The pipeline consists of three main stages: video segmentation, filtering, and automated bounding-box annotation.
Raw Video Trimming. Raw videos are trimmed into uniform three-second clips at 30 fps, discarding shorter segments to maintain temporal consistency.
Filtering. Video clips undergo automatic filtering based on person detection and scene-transition analysis. Specifically, we employed YOLOv11 [54] for human detection and PySceneDetect [55] to identify abrupt scene changes, discarding clips with insufficient human-action content or disrupted temporal continuity.
Bounding-box Annotation. Person bounding boxes are automatically annotated using YOLOv11 [54] detection combined with BoT-SORT tracking [56]. To enhance annotation accuracy, adaptive strategies were employed by assessing video motion intensity using sparse optical flow methods [57, 58, 59]. Annotation parameters were dynamically adjusted accordingly to optimize computational efficiency and annotation precision. The resulting structured annotations include bounding-box coordinates, unique identities, and detection confidence scores.
Manual Video Screening. Following automated preprocessing, an additional manual verification step was performed to ensure dataset purity and environmental consistency. Specifically, each generated video clip was individually reviewed to exclude terrestrial scenes, such as ground-based footage or pre-launch preparations, thereby retaining only those segments clearly depicting human activities under authentic microgravity conditions. This rigorous screening process ensured semantic precision and enhanced the overall reliability of the dataset.
Action Label Annotation. In this phase, we established a specialized action taxonomy tailored to microgravity environments, adapted from the AVA dataset’s original 80 atomic actions [12]. To ensure environmental applicability and semantic consistency, we systematically excluded irrelevant terrestrial or water-specific actions, merged closely related actions into unified categories, and introduced contextually appropriate semantic adjustments. Furthermore, clear differentiation criteria were defined for visually or semantically similar actions, thereby standardizing the annotation protocol. Subsequently, each three-second video clip was manually annotated as a discrete unit, assigning up to five distinct visible or inferable action labels per detected individual. The finalized annotations were exported as structured CSV files for subsequent utilization.
Caption and VQA Annotation. For caption annotation, we developed a rigorous protocol that combines detailed visual analysis of video content with supplementary contextual information sourced from official aerospace agency documentation to ensure both accuracy and authority. Specifically, we compiled and referenced mission objectives, crew rosters, and official reports from various space agencies across different countries and missions, enabling precise identification of individual astronauts. Furthermore, we utilized historical spacecraft layout diagrams and functional distribution maps of the spacecraft cabins to accurately determine the astronaut’s location within the spacecraft as well as the equipment inside it. In parallel, annotators conducted thorough reviews of the complete video sequences, identifying subtle differences between adjacent frames to precisely describe astronauts’ actions, physical appearances, and interactions with the environment. These high-quality captions provide strong semantic grounding to support downstream tasks such as action recognition and context-aware retrieval.
For Video Question Answering (VQA), we adopted a structured approach based on [60], formulating diverse questions using standard interrogative forms. Based on the complexity of the video captions, Large Language Models (LLMs) generate a proportional number of candidate question-answer pairs, explicitly filtering out cases with missing or uncertain information. These candidates are then ranked by the LLMs according to logical consistency, linguistic fluency, semantic relevance, and informational value. Subsequently, multiple rounds of human-assisted review and prompt adjustment are employed to eliminate any hallucinated or fabricated content produced by the LLMs. Finally, a concise set of six diverse and content-rich QA pairs is selected for each video, ensuring comprehensive coverage from broad contextual understanding to detailed actions.
Annotation Quality Control. During the annotation phase, a team of $9$ annotators collaboratively worked on labeling the fine-grained human actions, generating video captions, and crafting VQA pairs. To ensure high annotation quality, a comprehensive cross-verification protocol was implemented throughout the process, supported by group discussions to resolve disagreements and reach consensus. For the captioning and VQA tasks, all entries were further subjected to a semantic consistency check utilizing LLMs, followed by iterative human review to enhance both linguistic clarity and factual accuracy. This layered validation pipeline, combining automated and manual strategies, ensured the reliability and coherence of the annotations across all modalities. As a result, we introduce the MicroG-4M dataset, an extensively curated benchmark explicitly designed for fine-grained, multi-label, spatiotemporal action recognition, captioning, and VQA under microgravity conditions. This dataset sets a new foundation for advancing robust scene understanding and human activity analysis in space-based environments.
4 Dataset Composition
We release the MicroG-4M dataset, comprising the following subsets:
Fine-Grained Human Action Recognition Subset: Contains fine-grained, multi-label annotations for spatiotemporal human action detection, comprising $4,759$ manually annotated three-second video clips from authentic microgravity environments. Each clip includes annotations for up to five distinct actions per detected individual. The $50$ action labels are organized into three categories: Object Manipulation ($4,976$ annotations, $37.60\%$), Person Interaction ($1,198$ annotations, $32.34\%$), and Person Movement ($1,037$ annotations, $30.07\%$). Figure 2(d) illustrates the distribution of these categories in both real and simulated video clips, while Figure 2(e) shows the distribution of person counts per clip, highlighting the diversity of social configurations. Overall, the subset contains $13,261$ action annotations, including $9,610$ from real footage, $3,651$ from simulated sources, and $390,000$ bounding box annotations.
Video Caption Subset: Comprises $1,238$ detailed, semantically rich descriptions validated against official aerospace agency documentation for astronaut identities, spacecraft locations, actions, appearances, and interactions. Each caption corresponds uniquely to one video clip. The word cloud visualization in Figure 2(a) provides insights into frequently mentioned terms and concepts, illustrating the thematic and semantic distribution of the captions.
Visual Question Answering (VQA) Subset: Includes $7,428$ structured QA pairs, systematically generated and refined via LLMs to ensure linguistic fluency, semantic relevance, and comprehensive coverage of detailed actions and broader context, with each of the $1,238$ video clips associated with up to six diverse QA pairs. Figure 2(b) and (c) show word clouds for questions and answers separately, revealing prevalent inquiry types and common semantic patterns within the dataset.
Figure 2 presents representative video clips from the dataset, demonstrating typical annotation examples and highlighting the visual and semantic diversity of the dataset. More details regarding the dataset can be found in supplementary.
Figure 2: An illustration of the statistics of the dataset and the annotation samples. Word clouds of the (a) Caption, (b) Question, and (c) Answer from our MicroG-4M dataset are provided. The label statistics of the fine-grained human action recognition are provided in (d), which showcases the annotation number per action group (i.e., Object Manipulation (OM), Person Interaction (PI), Person Movement (PM)). The distribution of person counts per video clip is visualized in (e). On the bottom right, one annotation sample from MicroG-4M is provided.
5 Experiments Validating
5.1 Benchmark Protocol
Data Split. We partition the dataset into training, validation, and test subsets in a 7:1:2 ratio. Row-level distribution is as follows: Training set: 9,266 records (69.93% of 13,251); Validation set: 1,329 records (10.03% of 13,251); Test set: 2,656 records (20.04% of 13,251). Video-level distribution is as follows: Training set: 3,331 videos (69.99% of 4,759); Validation set: 475 videos (9.98% of 4,759); Test set: 953 videos (20.03% of 4,759).
Baselines for Fine-Grained HAR in Micro-Gravity Scenarios. For fine-grained action recognition, we evaluate well-established baselines from video-based human action recognition, including transformer-based models (MViTv1 [61], MViTv2 [16]) and CNN-based models (I3D [62], SlowFast [4], X3D [3], C2D [4]). These widely used architectures cover both paradigms and allow us to assess generalization to the unique spatiotemporal dynamics of microgravity. All models are initialized with Kinetics400 [11] pretrained weights for better convergence.
Baselines methods for Video Captioning and QA in Micro-Gravity Scenarios. For video captioning and question answering in microgravity, we evaluate strong baselines, including open-source models (VideoChatGPT [63], mPLUG-Owl-3 [64], LLaVA-Next [65], VideoLLaVa [66], Qwen-2.5-VL [67], InternVideo [17]) and closed-source models (ChatGPT-4o [19], Gemini 1.5 Pro [68]). Open-source models offer reproducibility via public weights and code, while closed-source models serve as upper-bound references. This mix enables both transparent analysis and comprehensive evaluation of video-language understanding in microgravity.
Evaluation Metrics for Fine-Grained HAR. Our evaluation metrics include mAP@0.5, F1 score, recall, and AUROC, all calculated using the macro method. Among these, mAP@0.5 is the primary metric for measuring average detection accuracy per category and thus comprehensively evaluating the model’s action recognition performance in a microgravity environment.
Evaluation Metrics for Video Caption and QA. We adopt standard automatic evaluation metrics, including CIDEr [69], BLEU-4 [70], ROUGE-L [71], METEOR [72], and BERTScore (F1) [73], all rescaled to a 0$\sim$100 range for consistency. For semantic similarity, we report S-BERT [74] and S-VQA [75] scores, both computed as cosine similarity between Sentence-BERT [76] embeddings of predicted and reference texts. S-VQA is used specifically for answer evaluation in generative visual question answering settings, capturing semantic equivalence beyond lexical overlap. More details regarding the implementations can be found in the supplementary materials.
5.2 Result Analysis for Fine-Grained Human Action Recognition
Quantitative Analysis.
Table 1: Performance comparison of models fine-tuned on MicroG-4M, evaluated on the validation and test sets.
Model Validation Test Arch TC Backbone #Params (M) mAP (%) F1-score (%) Recall (%) AUROC (%) mAP (%) F1-score (%) Recall (%) AUROC (%) C2D [4] 8x8 R50 [77] $23.61$ $27.22$ $12.52$ $10.34$ $82.86$ $29.51$ $8.09$ $6.58$ $83.49$ C2D NLN [4] 8x8 R50 [77] $30.97$ $40.42$ $23.10$ $20.41$ $87.11$ $44.64$ $28.30$ $24.86$ $89.40$ I3D [62] 8x8 R50 [77] $27.33$ $40.93$ $19.78$ $16.93$ $86.44$ $46.41$ $26.37$ $22.25$ $88.79$ I3D NLN [62] 8x8 R50 [77] $34.68$ $41.42$ $24.11$ $23.00$ $86.37$ $47.12$ $28.07$ $24.65$ $88.52$ Slow [4] 8x8 R50 [77] $31.74$ $40.32$ $21.83$ $19.08$ $84.55$ $45.19$ $26.13$ $22.77$ $88.49$ Slow [4] 4x16 R50 [77] $31.74$ $42.97$ $22.73$ $19.71$ $85.46$ $46.37$ $28.72$ $25.38$ $88.30$ SlowFast [4] 8x8 R50 [77] $33.76$ $38.76$ $20.29$ $17.66$ $85.91$ $43.02$ $22.63$ $18.98$ $88.51$ SlowFast [4] 4x16 R50 [77] $33.76$ $37.10$ $17.74$ $14.90$ $84.94$ $42.10$ $23.69$ $20.18$ $87.54$ MViTv1 [61] 16x4 B-CONV $36.34$ $17.79$ $7.89$ $6.86$ $72.40$ $12.86$ $5.54$ $4.66$ $74.63$ MViTv2 [16] 16x4 S $34.27$ $17.57$ $8.31$ $6.92$ $72.67$ $15.14$ $8.16$ $7.17$ $78.61$ X3D [3] 13x6 S $2.02$ $17.59$ $6.63$ $5.63$ $78.27$ $14.07$ $5.77$ $4.52$ $78.23$ X3D [3] 16x5 L $4.37$ $23.56$ $8.82$ $7.38$ $80.56$ $18.70$ $9.15$ $7.47$ $78.27$
Note: All models has been pretrained on Kinetics400 [11] dataset and continually trained on MicroG-4M. TC denotes the temporal configuration (frame length × sampling rate). #Params indicates the number of parameters (in millions, M).
Table 2: Zero-shot performance on MicroG-4M test set for models pretrained on Kinetics and fine-tuned on AVA [12].
Model Test Result Arch TC Backbone Pretrain Fine-tune mAP (%) F1-score (%) Recall (%) AUROC (%) Slow [4] 8x8 R50 [77] Kinetics 400 [11] AVA v2.2 [12] $16.24$ $2.67$ $1.99$ $73.83$ SlowFast [4] 32x2 R101 [77] Kinetics 600 [78] AVA v2.2 [12] $23.81$ $6.32$ $6.62$ $77.83$
Note: All metrics are macro-averaged over action classes. mAP is measured at IoU $=0.5$. F1 and AUROC are computed per class and then averaged. TC denotes the temporal configuration (frame length × sampling rate).
Figure 3: Qualitative results for fine-grained human action recognition in microgravity, where GT denotes ground truth, MicroG-4M indicates predictions from Slow [4] finetuned on MicroG-4M, and AVA denotes predictions from the same model finetuned on AVA [12]. The MicroG-4M model provides more accurate predictions than its Earth-trained counterpart.
We evaluate several representative models on MicroG-4M, all pretrained on Kinetics400 [11] and fine-tuned on our dataset. As shown in Table 1, most models plateau around $47\%$ test mAP, highlighting a substantial domain gap between Earth-based datasets and microgravity. The ranking of models deviates from trends on Kinetics400 [11] or AVA [12], e.g., convolutional models with non-local enhancements [4, 62] achieve the highest mAP and AUROC, while Slow (4×16) [4] yields the best F1-score, outperforming newer transformer-based designs [3, 16]. This suggests local spatial modeling and structured receptive fields remain advantageous in microgravity, where motion lacks gravitational consistency. Moreover, models with longer temporal windows perform better, reinforcing the importance of domain-adapted temporal reasoning for capturing slow or non-linear microgravity dynamics.
Comparing to AVA-finetuned models (Table 2), we find a substantial drop in performance when transferring directly from AVA [12] to MicroG-4M, even when pretraining and backbone are kept constant. For example, SlowFast trained on AVA [12] achieves only $23.81\%$ in mAP on MicroG-4M, significantly lower than its performance when fine-tuned on our dataset. This reinforces that Earth-trained models carry strong but fragile assumptions about vertical orientation, motion causality, and object affordances, which no longer hold in microgravity.
Together, these findings validate MicroG-4M’s role as a diagnostic benchmark that breaks gravity-dependent priors and challenges models to reason over irregular motion, spatial ambiguity, and contact-free interactions. Our dataset thus provides a rigorous testbed for evaluating and advancing space-adapted video understanding models, especially in the context of astronaut assistance and autonomous system development.
Qualitative Analysis. Figure 3 presents qualitative comparisons between models trained on AVA [12] and those fine-tuned on MicroG-4M, across 5 representative video clips captured inside and outside spacecraft cabins. The MicroG-4M–trained model demonstrates high alignment with ground truth labels for core actions such as “Stand”, “Walk”, and “Talk-to”, while the AVA [12]-based counterpart consistently misinterprets floating or inverted postures as “Bend/Bow” or “Sit”, revealing its reliance on Earth-centric gravitational priors.
A representative example in the fifth column shows the AVA [12]-finetuned model misclassifying a floating astronaut as “Sit” while the MicroG-4M–trained model correctly predicts “Stand” reflecting a common correction of gravity-induced biases. Models trained on MicroG-4M also demonstrate improved robustness in distinguishing passive object drift from intentional manipulation, though semantic ambiguity remains, e.g., predicting “Carry/Hold” when a tool drifts near the astronaut’s hand without actual interaction. These results indicate that MicroG-4M mitigates terrestrial biases, enhances sensitivity to body-object dynamics, and better captures domain-specific actions, supporting future work on temporal coherence and intentionality modeling.
5.3 Evaluation of Video Captioning Models
Table 3: Comparison of video captioning performance across open-source and closed-source models on the MicroG-4M benchmark. “#f” denotes the number of input frames used during model inference.
Model #f Cider Bleu-4 Rouge-L Meteor S-BERT BERTScore
Open Sourced
Video-ChatGPT [79] 3 0.06 0.12 10.10 4.33 39.61 85.40
mPLUG-Owl3 [64] 3 0.16 0.40 11.87 5.88 47.45 85.91
LLaVA-NeXT [80] 8 0.30 1.88 16.32 14.45 54.16 84.98
Video-LLaVA [66] 8 0.03 0.07 9.29 4.12 42.97 84.89
Qwen2.5-VL [81] 9 0.03 1.34 13.75 15.67 56.46 84.01
InternVideo [82] 90 0.77 2.60 16.57 15.18 55.28 85.41
Closed Sourced
GPT-4o [19] 6 1.74 2.65 16.46 11.27 62.18 86.75
Gemini 1.5 Pro [68] 16 3.52 3.28 17.34 15.19 63.38 86.25
The results in Table 3 reveal how video-language models perform on the MicroG-4M benchmark, highlighting key challenges introduced by microgravity-specific content. Lexical metrics such as CIDEr and BLEU-4 show low overall values, especially among open-source models, suggesting a significant distributional shift between MicroG-4M and typical pretraining data. The dataset’s domain-specific vocabulary, visually compositional scenes, and semantically dense annotations likely reduce surface-level overlap, which these metrics are sensitive to. In contrast, semantic similarity metrics such as S-BERT and BERTScore remain relatively higher and more consistent, indicating that several models capture the underlying intent even without lexical alignment. This underscores the semantic richness of MicroG-4M, where alternative phrasings and scientific terminology often convey similar meanings. Performance differences further reveal that input frame density and pretraining modality play key roles. InternVideo [17], which processes $90$ frames sampled within a $3$s window, consistently outperforms other open-source models. This suggests that dense sampling, coupled with video-specific pretraining, enhances the model’s ability to capture subtle spatial patterns and object-scene relationships—features that are particularly important in microgravity scenarios, where visual cues are often atypical or physically ambiguous. Closed-source models, i.e., GPT-4o [19] and Gemini 1.5 Pro [68] achieve better scores on both lexical and semantic metrics, likely due to broader data exposure, larger capacity, or more advanced cross-modal fusion strategies. However, their relatively small performance gains further validate the challenge posed by MicroG-4M. In general, these findings position MicroG-4M as a demanding benchmark for evaluating multimodal models under domain change, highlighting the need for robust spatial reasoning, domain adaptation, and semantically aware generation strategies in unconventional environments.
5.4 Evaluation of Visual Question Answering Models
Table 4: Performance comparison of Visual Question Answering (VQA) models on the MicroG-4M benchmark. “#f” denotes the number of input frames used during model inference.
Model #f Cider Bleu-4 Rouge-L Meteor S-VQA BERTScore
Open Sourced
LLaVA-NeXT [80] 8 24.00 22.14 15.56 12.40 38.08 87.15
Video-LLaVA [66] 8 25.70 28.47 15.90 10.71 35.39 87.13
Qwen2.5-VL [81] 9 2.99 0.65 8.35 8.47 40.65 84.80
Closed Sourced
Gemini 1.5 Pro [68] 16 8.78 1.33 13.03 12.54 43.15 86.41
GPT-4o [19] 6 33.98 3.76 18.11 15.89 44.56 87.81
The results in Table 4 demonstrate that MicroG-4M presents distinct challenges for visual question answering. Notably, there is a significant divergence between lexical and semantic evaluation metrics, particularly among open-source models. For example, Qwen2.5-VL [67] yields a BLEU-4 of only 0.65 and a CIDEr score of 2.99, yet achieves the highest S-VQA score in its category (40.65). This contrast suggests that MicroG-4M questions often admit multiple semantically valid answers that differ lexically, such as paraphrased actions, scientific terms, or object references adapted to microgravity settings. This characteristic does not reflect inconsistency in evaluation, but rather underscores the conceptual and linguistic diversity embedded in the dataset. In addition, the moderate absolute scores of even the top-performing closed-source models, such as GPT-4o [19] (CIDEr 33.98, S-VQA 44.56), reveal the difficulty of reasoning over visual content in microgravity. Unlike conventional VQA datasets, MicroG-4M includes visually ambiguous cues, e.g., floating objects, unusual body orientations, and tool manipulations under microgravity that challenge models trained primarily on terrestrial data. This suggests that current pretraining corpora lack sufficient coverage of such scenarios, and that purely scaling model capacity is insufficient for reliable generalization.
Interestingly, increasing the number of input frames within the fixed $3$s window does not consistently yield better performance. For example, Gemini 1.5 Pro [68] processes 16 frames but performs worse than GPT-4o [19], which uses only 6 frames. This indicates that dense frame sampling alone is insufficient. Instead, performance depends more critically on the model’s ability to extract semantically salient cues, e.g., astronaut posture, object manipulation, and spatial configurations, from visually subtle or low-motion segments. In microgravity environments, where conventional motion dynamics and object affordances are altered, effective spatial reasoning and cross-modal alignment appear to be more decisive than temporal redundancy. In summary, MicroG-4M reveals key limitations of current VQA systems in addressing domain-specific challenges, particularly those involving spatial complexity, ambiguous motion, and semantically flexible queries inherent to microgravity. Its comprehensive and specialized design establishes it as a valuable testbed for probing the robustness, adaptability, and generalization capabilities of multimodal models well beyond the scope of conventional Earth-based benchmarks.
6 Conclusion
In this work, we present MicroG-4M, the first large-scale dataset specifically curated for human action recognition and vision-language understanding in microgravity environments. The dataset features $4,759$ annotated video clips with over $390,000$ bounding boxes and $13,000$+ action labels across $50$ unique action classes. It also includes human-written captions and over $7,000$ VQA pairs, enabling rich semantic understanding and reasoning. To facilitate fair evaluation, we introduce MicroG-Bench, a benchmark that systematically assesses state-of-the-art models across fine-grained action recognition, video captioning, and question answering tasks. Our results show that existing deep learning models degrade significantly in space-like settings, emphasizing the need for dedicated benchmarks and domain-specific adaptation. MicroG-4M aims to advance robust, generalizable AI for astronaut support and autonomous space operations.
Limitations: The MicroG-4M dataset is currently limited to RGB input and includes both real and simulated footage, introducing potential modality and domain biases. Additionally, while the annotated subset is suitable for benchmarking, the scale of the dataset can be further enlarged to enable more space-related exploration for the human perception models. More illustrations can be found in the supplementary materials.
ACKNOWLEDGMENT
The project served to prepare the SFB 1574 Circular Factory for the Perpetual Product (project ID: 471687386), approved by the German Research Foundation (DFG, German Research Foundation). This work was supported in part by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026).