### v3

# Go Beyond Earth: Understanding Human Actions and Scenes in Microgravity Environments

Abstract
Despite substantial progress in video understanding, most existing datasets are limited to Earth’s gravitational conditions. However, microgravity alters human motion, interactions, and visual semantics, revealing a critical gap for real-world vision systems. This presents a challenge for domain-robust video understanding in safety-critical space applications. To address this, we introduce MicroG-4M, the first benchmark for spatio-temporal and semantic understanding of human activities in microgravity. Constructed from real-world space missions and cinematic simulations, the dataset includes 4,759 clips covering 50 actions, 1,238 context-rich captions, and over 7,000 question–answer pairs on astronaut activities and scene understanding. MicroG-4M aims to support three core tasks: fine-grained multi-label action recognition, temporal video captioning, and visual question answering, thereby enabling a comprehensive evaluation of both spatial localization and semantic reasoning in microgravity contexts. We establish baselines using state-of-the-art models. All data, annotations, and code are available at https://github.com/LEI-QI-233/HAR-in-Space.
1 Introduction
Yuri Gagarin’s historic flight in 1961 marked the beginning of human space exploration. Since then, significant milestones have been achieved, including crewed lunar landings, the continuous operation of the International Space Station (ISS) for over $25$ years, and the participation of more than $650$ individuals in space missions Moskowitz & Wolf (2025). With numerous planned crewed missions in the near future, the frequency and complexity of human activities in space are expected to increase substantially. In this context, ensuring the safety, enhancing the operational efficiency of space missions, and safeguarding the health and well-being of astronauts are of paramount importance.
With the rapid advancement of artificial intelligence, the integration of robotic systems aboard spacecraft is anticipated in the near future. These systems will assist astronauts in routine and mission-critical tasks. Consequently, there is a growing demand for the development of deep learning-based scene understanding and action recognition methods tailored specifically for the unique challenges posed by the microgravity environment.
Human action recognition Wang et al. (2023a); Feichtenhofer (2020); Feichtenhofer et al. (2019), scenario captioning Chen et al. (2024), and Visual Question Answering (VQA) Hu et al. (2023) are essential for intelligent human-robot collaboration, particularly in space, where precise perception and understanding of astronauts’ actions and their surrounding context are crucial for ensuring operational safety, efficiency, and autonomous assistance under constrained conditions. This capability is crucial for ensuring mission efficiency, enhancing astronaut safety, and providing autonomous assistance in the confined and complex conditions of space habitats. Numerous human action recognition datasets and video caption datasets have been developed, playing an important role in various research areas within computer vision and in industrial applications Camarena et al. (2023); Pareek & Thakkar (2021); Al-Faris et al. (2020); Le et al. (2022).
However, most of the existing datasets for video scenario captioning and action recognition are recorded on Earth without microgravity settings, e.g., Kinetics400 Kay et al. (2017), AVA Gu et al. (2018), FineGym Shao et al. (2020), and ActivityNet Captions Chen et al. (2019).
Human actions in space depart markedly from their terrestrial counterparts because microgravity removes gravity-aligned orientation and reliable support surfaces. Basic behaviors—standing, locomotion, eating, and manual manipulation—follow different kinematics and contact patterns Hagio et al. (2022). Standing becomes orientation-invariant, often maintained via foot restraints or handholds rather than ground support; locomotion is achieved by drifting or pulling along structures rather than gait; and manipulation frequently involves releasing or catching free-floating objects instead of placing or picking from a surface. These shifts violate terrestrial orientation, support/contact, and object-dynamics priors, which helps explain the degradation of Earth-trained Human Action Recognition (HAR) models in orbit and motivates a microgravity-specific benchmark.
Figure 1: An illustration of the MicroG-4M, containing videos from real and simulated microgravity environments (e.g., movies). The dataset supports benchmarks for three tasks: (1) video captioning, (2) video question answering, and (3) fine-grained human action recognition under microgravity.
To address this gap, we introduce MicroG-4M, a new video benchmark specifically designed for spatio-temporal and semantic understanding of human activities in microgravity. The name “4M” reflects four characteristics: Multi-source (Real mission footage and physically plausible film), Multimodal (RGB + text annotations), Multi-task (HAR, captioning, VQA), and Microgravity.
MicroG-4M comprises $4,759$ three-second video clips drawn from public YouTube footage of real space missions and carefully selected, realistic space-themed films to augment scenario diversity and coverage. The clips span scenes inside the ISS, the Tiangong Space Station, crewed spacecraft cabins, and extravehicular activities. The corpus contains more than $390,000$ annotated frames with bounding boxes for each visible individual and $13,000+$ action labels covering $50$ distinct actions. In addition, it includes $1,238$ descriptive clip captions created by human annotators and $7,000+$ open-ended question–answer pairs aimed at assessing factual, causal, and counterfactual understanding of scenes in microgravity. Both captions and QA annotations were carefully curated through multiple refinement rounds and data selection stages, ensuring high semantic fidelity, contextual accuracy, and relevance to the microgravity setting. Together, these resources enable multi-label spatio-temporal detection, fine-grained action recognition, caption generation, and VQA within a single dataset. To evaluate existing methods, we build the first comprehensive benchmark for these tasks, dubbed MicroG-Bench. We evaluate representative video encoders (SlowFast Feichtenhofer et al. (2019), X3D Feichtenhofer (2020), and MViT Li et al. (2022)) for human action recognition and leading vision-language models (e.g., InternVideo Wang et al. (2022), Gemini 1.5 Pro Anil et al. (2023), GPT-4o Hurst et al. (2024)) for captioning and VQA tasks. Across all evaluated subtasks, state-of-the-art terrestrial models experience significant performance degradation, underscoring the unique challenges posed by microgravity environments, such as arbitrary orientations, floating objects, and confined spacecraft interiors. These results highlight the necessity of dedicated benchmarks to facilitate the advancement of more robust and generalizable AI systems tailored for space applications. MicroG-Bench, therefore, provides a unified, fair yardstick that will guide the development of more robust and generalizable perception and language systems for astronaut assistance and autonomous mission operations. We summarize our contributions as follows. We present MicroG-4M, the first video-based dataset for activity recognition and scene understanding specifically designed for the microgravity environment. Additionally, we establish the first benchmark to evaluate state-of-the-art human action recognition methods and vision-language models for fine-grained action recognition and video captioning/QA in this unique setting.
2 Related Work
Vision-Based Understanding in Microgravity and Space Environments. Microgravity challenges terrestrial vision assumptions, requiring perception models for space habitats. Early work showed traditional SLAM to be unreliable, leading to robust alternatives with visual-inertial fusion, semantic mapping, CAD-informed constraints Soussan et al. (2022); Mao et al. (2024); Miller et al. (2022); Tweddle et al. (2015), and validated on platforms like Astrobee Kang et al. (2024). Real-time vision has also enabled dynamic tasks such as target tracking and scene change detection Oestreich et al. (2021); Dinkel et al. (2024). Research has shifted toward interaction, including astronaut pose recovery Gan et al. (2023); Ouyang et al. (2025), gesture recognition Lingyun et al. (2020); Gao et al. (2020), and EMG-based input Assad et al. (2013). Yet, high-level semantic understanding of astronaut actions and intent remains limited. Current assistants like CIMON Eisenberg et al. (2024) provide basic perception but lack contextual reasoning. To address this, we propose a unified benchmark for action recognition, video captioning, and visual question answering in space operations.
Datasets for Action Detection, Video Captioning, and VQA. Progress in video understanding has been largely enabled by the introduction of benchmark datasets across action detection, captioning, and VQA. For action detection, early datasets such as UCF101 Soomro et al. (2012) and HMDB51 Kuehne et al. (2011) provided trimmed classification tasks, later extended by Kinetics Kay et al. (2017) and ActivityNet Caba Heilbron et al. (2015) to large-scale, untrimmed settings with diverse action categories. AVA Gu et al. (2018) further introduced spatio-temporal annotations for atomic actions, enabling fine-grained multi-label detection in realistic scenes. Video captioning evolved from MSVD Chen & Dolan (2011) and MSR-VTT Xu et al. (2016), which paired short clips with multiple sentences, to dense captioning in untrimmed videos via ActivityNet Captions Krishna et al. (2017) and procedural datasets like YouCook2 Zhou et al. (2018). Multilingual benchmarks such as VATEX Wang et al. (2019) expanded the task to cross-lingual settings with high-quality parallel annotations. In VQA, static image datasets like VQA v2.0 Goyal et al. (2017) and CLEVR Johnson et al. (2017) laid the foundation for compositional reasoning, while TGIF-QA Jang et al. (2017), MovieQA Tapaswi et al. (2016), and TVQA Lei et al. (2018) introduced temporal and multimodal reasoning in video. Recent efforts such as NExT-QA Xiao et al. (2021) and CLEVRER Yi et al. (2020) focus on causal inference and counterfactual reasoning. However, existing datasets are constrained to terrestrial environments and lack domain-specific complexities inherent to space-based activities. To bridge this gap, we introduce the first benchmark for video captioning and VQA in microgravity, enabling the evaluation of vision-language models in long-horizon, space-relevant scenarios.
Fine-Grained Human Action Recognition. Fine-grained video-based action recognition Gritsenko et al. (2023); Chung et al. (2021) aims to detect fine-grained, indivisible human actions in both single- and multi-person videos. Unlike standard clip-level recognition, atomic action localization requires frame-level, multi-label classification with spatio-temporal bounding box predictions. To tackle this, CNN-based Wang et al. (2023a); Feichtenhofer (2020); Feichtenhofer et al. (2019) and transformer-based Ryali et al. (2023); Wang et al. (2023a); Li et al. (2022); Wang et al. (2023b; 2022); Peng et al. (2022); Gritsenko et al. (2023) models have been adapted by adding multi-label heads, bounding box regressors, and region-of-interest modules. Notably, Ryali et al.Ryali et al. (2023) introduced a hierarchical vision transformer balancing accuracy and efficiency, while Wanget al. Wang et al. (2023a) proposed a dual-masked autoencoder for improved video pretraining. In this work, we evaluate such HAR methods for the first time in microgravity scenarios.
3 Collection Methodology
We aim to construct a dataset comprising authentic footage recorded aboard the International Space Station and other spacecraft, as well as films that realistically depict microgravity conditions. In the following, we introduce our comprehensive pipeline for collecting, assembling, filtering, and annotating the MicroG-4M dataset, which supports multi-label spatio-temporal action detection, video captioning, and video question answering tasks from Internet-sourced videos.
Raw Video Information Collection. Video sources primarily include genuine microgravity footage from actual spacecraft missions and selected cinematic clips known for their realistic depictions of weightlessness. Authentic spacecraft videos were mainly retrieved from online video platforms, while cinematic clips were manually chosen based on their fidelity to real microgravity conditions. All videos were standardized to a resolution of $480p$ to maintain consistency. The final dataset comprises approximately $5,000$ three-second clips, predominantly composed of authentic spacecraft footage.
Dataset Assembly Pipeline. We established an automated pipeline to preprocess and structure raw videos, enabling consistent and accurate downstream annotation. The pipeline consists of three main stages: video segmentation, filtering, and automated bounding-box annotation.
Raw Video Trimming. Raw videos are trimmed into uniform three-second clips at 30 fps, discarding shorter segments to maintain temporal consistency.
Filtering. Video clips undergo automatic filtering based on person detection and scene-transition analysis. Specifically, we employed YOLOv11 Khanam & Hussain (2024) for human detection and PySceneDetect Castellano to identify abrupt scene changes, discarding clips with insufficient human-action content or disrupted temporal continuity.
Bounding-box Annotation. Person bounding boxes are automatically annotated using YOLOv11 Khanam & Hussain (2024) detection combined with BoT-SORT tracking Aharon et al. (2022). To enhance annotation accuracy, adaptive strategies were employed by assessing video motion intensity using sparse optical flow methods Jeannin & Divakaran (2001); Ali (2013); Szeliski (2010). Annotation parameters were dynamically adjusted accordingly to optimize computational efficiency and annotation precision. The resulting structured annotations include bounding-box coordinates, unique identities, and detection confidence scores.
Manual Video Screening. Following automated preprocessing, an additional manual verification step was performed to ensure dataset purity and environmental consistency. Specifically, each generated video clip was individually reviewed to exclude terrestrial scenes, such as ground-based footage or pre-launch preparations, thereby retaining only those segments clearly depicting human activities under authentic microgravity conditions. This rigorous screening process ensured semantic precision and enhanced the overall reliability of the dataset.
[…]
Caption and VQA Annotation. For caption annotation, we developed a rigorous protocol that combines detailed visual analysis of video content with supplementary contextual information sourced from official aerospace agency documentation to ensure both accuracy and authority. Specifically, we compiled and referenced mission objectives, crew rosters, and official reports from various space agencies across different countries and missions, enabling precise identification of individual astronauts. Furthermore, we utilized historical spacecraft layout diagrams and functional distribution maps of the spacecraft cabins to accurately determine the astronaut’s location within the spacecraft as well as the equipment inside it. In parallel, annotators conducted thorough reviews of the complete video sequences, identifying subtle differences between adjacent frames to precisely describe astronauts’ actions, physical appearances, and interactions with the environment. These high-quality captions provide strong semantic grounding to support downstream tasks such as action recognition and context-aware retrieval.
For Video Question Answering (VQA), we adopted a structured approach based on Heilman & Smith (2009), formulating diverse questions using standard interrogative forms. Based on the complexity of the video captions, Large Language Models (LLMs) generate a proportional number of candidate question-answer pairs, explicitly filtering out cases with missing or uncertain information. These candidates are then ranked by the LLMs according to logical consistency, linguistic fluency, semantic relevance, and informational value. Subsequently, multiple rounds of human-assisted review and prompt adjustment are employed to eliminate any hallucinated or fabricated content produced by the LLMs. Finally, a concise set of six diverse and content-rich QA pairs is selected for each video, ensuring comprehensive coverage from broad contextual understanding to detailed actions.
Annotation Quality Control. During the annotation phase, a team of $9$ annotators collaboratively worked on labeling the fine-grained human actions, generating video captions, and crafting VQA pairs. To ensure high annotation quality, a comprehensive cross-verification protocol was implemented throughout the process, supported by group discussions to resolve disagreements and reach consensus. For the captioning and VQA tasks, all entries were further subjected to a semantic consistency check utilizing LLMs, followed by iterative human review to enhance both linguistic clarity and factual accuracy. This layered validation pipeline, combining automated and manual strategies, ensured the reliability and coherence of the annotations across all modalities. As a result, we introduce the MicroG-4M dataset, an extensively curated benchmark explicitly designed for fine-grained, multi-label, spatiotemporal action recognition, captioning, and VQA under microgravity conditions. This dataset sets a new foundation for advancing robust scene understanding and human activity analysis in space-based environments.
4 Dataset Composition
We release the MicroG-4M dataset, comprising the following subsets:
Figure 2: An illustration of the statistics of the dataset and the annotation samples. Word clouds of the (a) Caption, (b) Question, and (c) Answer from our MicroG-4M dataset are provided. The label statistics of the fine-grained human action recognition are provided in (d), which showcases the annotation number per action group (i.e., Object Manipulation (OM), Person Interaction (PI), Person Movement (PM)). The distribution of person counts per video clip is visualized in (e). On the bottom right, one annotation sample from MicroG-4M is provided.
Fine-Grained Human Action Recognition Subset: Contains fine-grained, multi-label annotations for spatiotemporal human action detection, comprising $4,759$ manually annotated three-second video clips from authentic microgravity environments. Each clip includes annotations for up to five distinct actions per detected individual. The $50$ action labels are organized into three categories: Object Manipulation ($4,976$ annotations, $37.60\%$), Person Interaction ($4,288$ annotations, $32.34\%$), and Person Movement ($3,987$ annotations, $30.07\%$). Figure 2(d) illustrates the distribution of these categories in both real and simulated video clips, while Figure 2(e) shows the distribution of person counts per clip, highlighting the diversity of social configurations. Overall, the subset contains $13,261$ action annotations, including $9,610$ from real footage, $3,651$ from simulated sources, and $390,000$ bounding box annotations.
Video Caption Subset: Comprises $1,238$ detailed, semantically rich descriptions validated against official aerospace agency documentation for astronaut identities, spacecraft locations, actions, appearances, and interactions. Each caption corresponds uniquely to one video clip. The word cloud visualization in Figure 2(a) provides insights into frequently mentioned terms and concepts, illustrating the thematic and semantic distribution of the captions.
Visual Question Answering (VQA) Subset: Includes $7,428$ structured QA pairs, systematically generated and refined via LLMs to ensure linguistic fluency, semantic relevance, and comprehensive coverage of detailed actions and broader context, with each of the $1,238$ video clips associated with up to six diverse QA pairs. Figure 2(b) and (c) show word clouds for questions and answers separately, revealing prevalent inquiry types and common semantic patterns within the dataset.
Figure 2 presents representative video clips from the dataset, demonstrating typical annotation examples and highlighting the visual and semantic diversity of the dataset. More details regarding the dataset can be found in the appendix.
5 Experiments Validating
5.1 Benchmark Protocol
Data Split. We partition the dataset into training, validation, and test subsets in a 7:1:2 ratio. Row‐level distribution is as follows: Training set: 9,266 records (69.93% of 13,251); Validation set: 1,329 records (10.03% of 13,251); Test set: 2,656 records (20.04% of 13,251). Video‐level distribution is as follows: Training set: 3,331 videos (69.99% of 4,759); Validation set: 475 videos (9.98% of 4,759); Test set: 953 videos (20.03% of 4,759).
Baselines for Fine-Grained HAR in Microgravity Scenarios. For fine-grained action recognition, we evaluate well-established baselines from video-based human action recognition, including transformer-based models (MViTv1 Fan et al. (2021), MViTv2 Li et al. (2022)) and CNN-based models (I3D Carreira & Zisserman (2017), SlowFast Feichtenhofer et al. (2019), X3D Feichtenhofer (2020), C2D Feichtenhofer et al. (2019)). These widely used architectures cover both paradigms and allow us to assess generalization to the unique spatiotemporal dynamics of microgravity. All models are initialized with Kinetics400 Kay et al. (2017) pretrained weights for better convergence.
Baseline methods for Video Captioning and QA in Microgravity Scenarios. For video captioning and question answering in microgravity, we evaluate strong baselines, including open-source models (VideoChatGPT Li et al. (2023), mPLUG-Owl-3 Ye et al. (2025), LLaVA-Next Li et al. (2024), VideoLLaVa Lin et al. (2024), Qwen-2.5-VL Bai et al. (2025), InternVideo Wang et al. (2022)) and closed-source models (GPT-4o Hurst et al. (2024), Gemini 1.5 Pro Reid et al. (2024)). Open-source models offer reproducibility via public weights and code, while closed-source models serve as upper-bound references. This mix enables both transparent analysis and comprehensive evaluation of video-language understanding in microgravity.
Cross-domain transfer protocol. To quantify the microgravity-induced domain gap, we use a fixed transfer setup: models pretrained on Kinetics are fine-tuned on AVA Gu et al. (2018) with matched settings, then evaluated zero-shot on MicroG-4M. For terrestrial contrast, we test on JHMDB (Split 1) Jhuang et al. (2013) with the overlapping action set and standard protocol. This isolates domain effects (microgravity vs. Earth) from implementation choices; evaluation details in Sec. 5.
Evaluation Metrics for Fine-Grained HAR. Our evaluation metrics include mAP@0.5, F1 score, recall, and AUROC, all calculated using the macro method. Among these, mAP@0.5 is the primary metric for measuring average detection accuracy per category and thus comprehensively evaluating the model’s action recognition performance in a microgravity environment. Per-class threshold sweeps are provided in the appendix.
[…]
5.2 Result Analysis for Fine-Grained Human Action Recognition
[…]
Quantitative Analysis.
Table 1: Performance of models fine-tuned on MicroG-4M, evaluated on the validation and test sets.
[…]
Note: All models have been pretrained on Kinetics400 Kay et al. (2017) dataset and continually trained on MicroG-4M. TC denotes the temporal configuration (frame length × sampling rate). #Params indicates the number of parameters (in millions, M).
Table 2: Zero-shot performance on MicroG-4M test set for models pretrained on Kinetics and fine-tuned on AVA Gu et al. (2018).
Model Test Result Arch TC Backbone Pretrain Fine-tune mAP (%) F1-score (%) Recall (%) AUROC (%) Slow Feichtenhofer et al. (2019) 8x8 R50 He et al. (2016) Kinetics 400 Kay et al. (2017) AVA v2.2 Gu et al. (2018) $16.24$ $2.67$ $1.99$ $73.83$ SlowFast Feichtenhofer et al. (2019) 32x2 R101 He et al. (2016) Kinetics 600 Carreira et al. (2018) AVA v2.2 Gu et al. (2018) $23.81$ $6.32$ $6.62$ $77.83$
Note: All metrics are macro-averaged over action classes. mAP is measured at IoU =0.5=0.5. F1 and AUROC are computed per class and then averaged. TC denotes the temporal configuration (frame length × sampling rate).
Table 3: Cross-domain transfer under matched AVA fine-tuning (zero-shot evaluation).
Model TC Backbone Test Set mAP (%) AUROC (%)
SlowFast Feichtenhofer et al. (2019) $32{\times}2$ R101 He et al. (2016) JHMDB Jhuang et al. (2013) 47.50 83.98
SlowFast Feichtenhofer et al. (2019) $32{\times}2$ R101 He et al. (2016) MicroG-4M 23.81 77.83
Slow Feichtenhofer et al. (2019) $8{\times}8$ R50 He et al. (2016) JHMDB Jhuang et al. (2013) 34.24 76.96
Slow Feichtenhofer et al. (2019) $8{\times}8$ R50 He et al. (2016) MicroG-4M 16.24 73.83
Figure 3: Qualitative results for fine-grained human action recognition in microgravity, where GT denotes ground truth, MicroG-4M indicates predictions from Slow fine-tuned on MicroG-4M, and AVA denotes predictions from the same model fine-tuned on AVA. The MicroG-4M model provides more accurate predictions than its Earth-trained counterpart.
[…]
Qualitative Analysis. Figure 3 presents qualitative comparisons between models trained on AVA Gu et al. (2018) and those fine-tuned on MicroG-4M, across 5 representative video clips captured inside and outside spacecraft cabins. The MicroG-4M–trained model demonstrates high alignment with ground truth labels for core actions such as “Stand”, “Walk”, and “Talk-to”, while the AVA Gu et al. (2018)-based counterpart consistently misinterprets floating or inverted postures as “Bend/Bow” or “Sit”, revealing its reliance on Earth-centric gravitational priors. A representative example in the fifth column shows the AVA Gu et al. (2018)-finetuned model misclassifying a floating astronaut as “Sit” while the MicroG-4M–trained model correctly predicts “Stand” reflecting a common correction of gravity-induced biases. Models trained on MicroG-4M also demonstrate improved robustness in distinguishing passive object drift from intentional manipulation, though semantic ambiguity remains, e.g., predicting “Carry/Hold” when a tool drifts near the astronaut’s hand without actual interaction. These results indicate that MicroG-4M mitigates terrestrial biases, enhances sensitivity to body-object dynamics, and better captures domain-specific actions, supporting future work on temporal coherence and intentionality modeling.
5.3 Evaluation of Video Captioning Models
Table 4: Comparison of video captioning performance across open-source and closed-source models on the MicroG-4M benchmark. “#f” denotes the number of input frames used during model inference.
Model #f CIDEr BLEU-4 Rouge-L Meteor S-BERT BERTScore
[…]
Video-ChatGPT Maaz et al. (2024) 3 0.06 0.12 10.10 4.33 39.61 85.40
mPLUG-Owl3 Ye et al. (2025) 3 0.16 0.40 11.87 5.88 47.45 85.91
LLaVA-NeXT Zhang et al. (2024) 8 0.30 1.88 16.32 14.45 54.16 84.98
Video-LLaVA Lin et al. (2024) 8 0.03 0.07 9.29 4.12 42.97 84.89
Qwen2.5-VL Qwen et al. (2025) 9 0.03 1.34 13.75 15.67 56.46 84.01
Tarsier2-Recap-7B Yuan et al. (2025) 16 0.04 0.03 0.17 0.12 51.35 84.53
InternVideo Xing et al. (2024) 90 0.77 2.60 16.57 15.18 55.28 85.41
[…]
GPT-4o Hurst et al. (2024) 6 1.74 2.65 16.46 11.27 62.18 86.75
Gemini 1.5 Pro Reid et al. (2024) 16 3.52 3.28 17.34 15.19 63.38 86.25
The results in Table 5.3 reveal how video-language models perform on the MicroG-4M benchmark, highlighting key challenges introduced by microgravity-specific content. Lexical metrics such as CIDEr and BLEU-4 show low overall values, especially among open-source models, suggesting a significant distributional shift between MicroG-4M and typical pretraining data. The dataset’s domain-specific vocabulary, visually compositional scenes, and semantically dense annotations likely reduce surface-level overlap, which these metrics are sensitive to. In contrast, semantic similarity metrics such as S-BERT and BERTScore remain relatively higher and more consistent, indicating that several models capture the underlying intent even without lexical alignment. This underscores the semantic richness of MicroG-4M, where alternative phrasings and scientific terminology often convey similar meanings. Performance differences further reveal that input frame density and pretraining modality play key roles. InternVideo Wang et al. (2022), which processes $90$ frames sampled within a $3$s window, consistently outperforms other open-source models. This suggests that dense sampling, coupled with video-specific pretraining, enhances the model’s ability to capture subtle spatial patterns and object-scene relationships—features that are particularly important in microgravity scenarios, where visual cues are often atypical or physically ambiguous. Closed-source models, i.e., GPT-4o Hurst et al. (2024) and Gemini 1.5 Pro Reid et al. (2024) achieve better scores on both lexical and semantic metrics, likely due to broader data exposure, larger capacity, or more advanced cross-modal fusion strategies. However, their relatively small performance gains further validate the challenge posed by MicroG-4M. In general, these findings position MicroG-4M as a demanding benchmark for evaluating multimodal models under domain change, highlighting the need for robust spatial reasoning, domain adaptation, and semantically aware generation strategies in unconventional environments.
5.4 Evaluation of Visual Question Answering Models
Table 5: Experiments of Visual Question Answering (VQA) models on the MicroG-4M benchmark. “#f” denotes the number of input frames used during model inference.
Model #f CIDEr BLEU-4 Rouge-L Meteor S-VQA BERTScore
[…]
LLaVA-NeXT Zhang et al. (2024) 8 24.00 22.14 15.56 12.40 38.08 87.15
Video-LLaVA Lin et al. (2024) 8 25.70 28.47 15.90 10.71 35.39 87.13
Qwen2.5-VL Qwen et al. (2025) 9 2.99 0.65 8.35 8.47 40.65 84.80
Tarsier2-Recap-7B Yuan et al. (2025) 16 5.08 0.01 0.08 0.09 29.60 85.30
[…]
Gemini 1.5 Pro Reid et al. (2024) 16 8.78 1.33 13.03 12.54 43.15 86.41
GPT-4o Hurst et al. (2024) 6 33.98 3.76 18.11 15.89 44.56 87.81
The results in Table 5.4 demonstrate that MicroG-4M presents distinct challenges for visual question answering. Notably, there is a significant divergence between lexical and semantic evaluation metrics, particularly among open-source models. For example, Qwen2.5-VL Bai et al. (2025) yields a BLEU-4 of only 0.65 and a CIDEr score of 2.99, yet achieves the highest S-VQA score in its category (40.65). This contrast suggests that MicroG-4M questions often admit multiple semantically valid answers that differ lexically, such as paraphrased actions, scientific terms, or object references adapted to microgravity settings. This characteristic does not reflect inconsistency in evaluation, but rather underscores the conceptual and linguistic diversity embedded in the dataset. In addition, the moderate absolute scores of even the top-performing closed-source models, such as GPT-4o Hurst et al. (2024) (CIDEr 33.98, S-VQA 44.56), reveal the difficulty of reasoning over visual content in microgravity. Unlike conventional VQA datasets, MicroG-4M includes visually ambiguous cues, e.g., floating objects, unusual body orientations, and tool manipulations under microgravity that challenge models trained primarily on terrestrial data. This suggests that current pretraining corpora lack sufficient coverage of such scenarios, and that purely scaling model capacity is insufficient for reliable generalization. Interestingly, increasing the number of input frames within the fixed $3$s window does not consistently yield better performance. For example, Gemini 1.5 Pro Reid et al. (2024) processes 16 frames but performs worse than GPT-4o Hurst et al. (2024), which uses only 6 frames. This indicates that dense frame sampling alone is insufficient. Instead, performance depends more critically on the model’s ability to extract semantically salient cues, e.g., astronaut posture, object manipulation, and spatial configurations, from visually subtle or low-motion segments. In microgravity environments, where conventional motion dynamics and object affordances are altered, effective spatial reasoning and cross-modal alignment appear to be more decisive than temporal redundancy. In summary, MicroG-4M reveals key limitations of current VQA systems in addressing domain-specific challenges, particularly those involving spatial complexity, ambiguous motion, and semantically flexible queries inherent to microgravity. Its comprehensive and specialized design establishes it as a valuable testbed for probing the robustness, adaptability, and generalization capabilities of multimodal models well beyond the scope of conventional Earth-based benchmarks.
6 Conclusion
In this work, we present MicroG-4M, the first large-scale dataset specifically curated for human action recognition and vision-language understanding in microgravity environments. The dataset features $4,759$ annotated video clips with over $390,000$ bounding boxes and 13,00013,000+ action labels across $50$ unique action classes. It also includes human-written captions and over $7,000$ VQA pairs, enabling rich semantic understanding and reasoning. We introduce MicroG-Bench, a benchmark for evaluating state-of-the-art models in fine-grained action recognition, video captioning, and question answering. Results show significant performance degradation in space-like settings, highlighting the need for domain-specific benchmarks and adaptation. MicroG-4M advances robust, generalizable AI for astronaut support and autonomous space operations.
Acknowledgement
The project served to prepare the SFB 1574 Circular Factory for the Perpetual Product (project ID: 471687386), approved by the German Research Foundation (DFG, German Research Foundation). This work was supported in part by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026). This project is also supported in part by the National Natural Science Foundation of China under Grant No. 62473139, in part by the Hunan Provincial Research and Development Project (Grant No. 2025QK3019), and in part by the Open Research Project of the State Key Laboratory of Industrial Control Technology, China (Grant No. ICT2025B20).
Appendix
Appendix A Technical Limitations
MicroG-4M introduces the first unified benchmark for high-level video understanding in microgravity environments, but several technical limitations remain that could guide future refinement and expansion.
One challenge lies in the ambiguity and subjectivity of annotations. Interpreting actions and generating captions can vary between annotators, especially in visually ambiguous frames or when inference is required. Even with cross-validation protocols in place, some degree of subjectivity may persist, introducing annotation noise.
Another limitation is the restricted temporal context. Each video clip spans only three seconds, which may hinder the ability to model long-term dependencies. This limitation is particularly relevant when modeling multi-step operations that are common in space missions.
The dataset is currently limited to RGB visual input, which constrains the potential for multimodal understanding. Future versions could benefit from the inclusion of additional modalities such as audio signals or communication transcripts to support more comprehensive reasoning.
There is also an inherent domain bias introduced by the inclusion of cinematic footage. While these clips are visually high-fidelity and physically plausible, they may differ in visual style and narrative framing from real operational recordings. This discrepancy can affect the generalizability of models trained on the dataset.
Finally, the scale and annotation coverage of the dataset present further constraints. The current release offers a carefully annotated subset of the full video collection, suitable for benchmarking, but smaller than many large-scale web datasets. Continued annotation efforts are underway to expand the dataset, covering a wider range of clips, actions, and scene types to support more diverse downstream applications, including temporal reasoning and sequence-level inference.
Appendix B Potential Social Impacts
MicroG-4M introduces the first benchmark specifically designed for video understanding and vision-language reasoning in microgravity environments. It comprises 4,759 curated clips supporting fine-grained action recognition, video captioning, and visual question answering. These tasks collectively provide a testbed for evaluating models under the unique motion dynamics and spatial ambiguities posed by microgravity.
While MicroG-4M is not intended for deployment, it may inform downstream research in areas such as astronaut behavior modeling, procedural understanding, or human–robot collaboration. Its vision-language annotations also facilitate studies on video summarization and temporal grounding, with potential implications for future human-AI interfaces in space-based or analog settings.
MicroG-4M enables the analysis of model limitations under microgravity-specific challenges, such as sensitivity to gravitational priors and orientation ambiguity, offering a foundation for research on robustness and generalization. Although developed for space-based contexts, the dataset’s motion and interaction patterns may inspire comparative studies in gravity-reduced analog environments, such as underwater settings.
Furthermore, the semantic complexity of the captioning and QA tasks highlights challenges including hallucination and semantic inconsistency, positioning MicroG-4M as a testbed for evaluating the reliability and grounding of multimodal models in physically unfamiliar conditions.
Appendix C Safety and Ethical Discussion
MicroG-4M is developed as a research-oriented benchmark for video understanding in microgravity environments. While ethical standards were followed throughout its construction, several considerations are noted to promote responsible and informed use.
All real-world videos are sourced from publicly available materials released by official space agencies and educational institutions, containing no private or sensitive content. Astronauts are shown exclusively in professional contexts. Simulated cinematic content, while enriching visual diversity, may introduce stylistic bias that differs from real operational footage.
All captions and QA annotations were created manually by annotators with domain guidance. LLMs were employed exclusively for grammatical correction and fluency enhancement, without contributing to semantic generation. Nonetheless, users should remain aware of any residual stylistic biases introduced during this refinement process.
MicroG-4M is intended solely for non-commercial academic research. It is not validated for real-world deployment, particularly in sensitive domains such as surveillance or defense. Users are advised to evaluate generalization carefully and avoid overextension of model outputs.
We encourage community feedback on potential biases, content issues, or safety risks. Future versions will include expanded validation and filtering to enhance transparency and data quality.
Appendix D Additional Qualitative Analysis
D.1 Fine-Grained Human Action Recognition
We provide qualitative comparisons using representative video samples to illustrate the fine-grained action recognition performance of different models in microgravity scenarios (Figure 4).
The AVA fine-tuned model frequently misinterprets microgravity-specific postures and contexts, exhibiting errors such as misclassifying hatch-crossing as bend/bow or generating unrealistic predictions like detecting a “smoke” action. In contrast, the MicroG-4M fine-tuned models demonstrate improved accuracy by correctly identifying nuanced and compound actions specific to microgravity environments, as illustrated by the more precise recognition of actions like carry/hold, lift/pickup, and put on/off clothing. The highest-performing MicroG-4M model (I3D Non-local) further reduces both false positives and missed detections, highlighting its superior capability in recognizing complex and subtle astronaut activities.
These examples underscore the necessity for specialized training data, such as MicroG-4M, to effectively capture and recognize fine-grained human actions unique to microgravity conditions.
Figure 4: Qualitative results for fine-grained human action recognition in microgravity. Below the frame samples, the first row presents the ground truth labels of the actions. The second row presents the predictions of the Slow architecture fine-tuned on the AVA dataset. The third row shows the predictions of the Slow architecture fine-tuned on the MicroG-4M dataset. The last row shows the predictions of the I3D Non-Local Network (NLN) architecture fine-tuned on MicroG-4M. For both the Slow and I3D NLN architectures, the AVA and MicroG-4M models were trained under the same configuration: a ResNet-50 backbone with an 8×8 input (frame length × sampling rate), pre-trained on Kinetics-400. The I3D NLN model fine-tuned on MicroG-4M achieved the highest mAP among our baselines. Gray text denotes missed detections, while red text denotes false detections.
D.2 Video Captioning
We provide qualitative examples to emphasize the unique challenges posed by space-related scenarios and the limitations of existing state-of-the-art multimodal models (mPLUG-Owl3, LLaVA-NeXT, GPT-4o, Gemini 1.5 Pro) in accurately capturing specialized details inherent to space station environments.
In the first scenario (Figure 5, left), all models demonstrate significant shortcomings in capturing critical, astronaut-specific information and precise operational context, underscoring the difficulty in accurately describing space-station activities without access to specialized annotations.
The second scenario (Figure 5, right) represents a relatively simpler context, yet models still lack precise details and fail to fully leverage the specialized information present in our ground-truth annotations. These examples illustrate the inherent difficulty in accurately modeling highly specialized and contextually rich scenarios typical of space environments, highlighting the necessity and distinctive value of our carefully annotated space-oriented dataset.
Overall, these qualitative results demonstrate the critical importance of domain-specific annotation for effectively capturing the nuanced details and specialized context essential in space exploration scenarios.
Figure 5: Qualitative examples illustrating ground-truth captions and outputs from four state-of-the-art multimodal models. The left example represents a challenging scenario, in which all models fail to accurately capture detailed and precise information. The right example demonstrates a relatively simpler scenario, where model-generated captions exhibit closer alignment to the ground truth. Each caption includes the corresponding caption length (in words), with key details highlighted in the ground-truth captions.
D.3 Video Question Answering
We present qualitative examples illustrating the performance of state-of-the-art multimodal models (Gemini 1.5 Pro, GPT-4o, Video-LLaVA) on challenging Video Question-Answering (VQA) tasks specifically related to space environments.
The provided examples (Figure 6) showcase diverse and challenging scenarios uniquely associated with space-based contexts. The top-left example demonstrates the complexity of accurately interpreting multiple astronauts’ actions within confined spaces, especially when occlusions occur. The bottom-left example emphasizes challenges in accurately identifying specific locations unique to space environments. The top-right scenario tests models’ abilities to describe various specialized objects within densely detailed backgrounds, typical of spacecraft interiors. Lastly, the bottom-right example addresses the models’ propensity for hallucination by evaluating their capacity to accurately infer details not explicitly mentioned in the visual content.
Overall, these qualitative analyses highlight significant limitations of current multimodal models in handling space-specific scenarios, underscoring the critical need for detailed and specialized annotations provided by our dataset to enhance performance and robustness in space-related VQA tasks.
Figure 6: Qualitative examples illustrating Ground-Truth (GT) answers and corresponding responses generated by three state-of-the-art multimodal models.
Appendix E Data Format Conversion
To simplify frame-level annotation processing, we adapted our dataset to match the AVA format, with one modification: using frame stamps rather than timestamps. This allows direct frame indexing, aligning better with our annotation scheme and reducing unnecessary conversion overhead.
E.1 Dataset Metadata
Our dataset metadata is structured hierarchically into two categories—video-level and frame-level—to facilitate efficient retrieval and flexible querying.
Video-Level Metadata: summarizes high-level annotations and semantic context:
  • Video ID: Unique identifier for each video clip.
  • Person ID: Unique identifier for annotated individuals within videos.
  • Action Labels: Fine-grained human action labels.
  • Caption: Detailed textual description of the video content.
  • Question-Answer Pairs: Structured questions and answers related to video content.
Frame-Level Metadata: provides detailed annotations at individual frame granularity:
  • Video ID: Corresponding identifier linking video-level metadata.
  • Frame stamp: Temporal location of annotated frames within videos.
  • Person ID: Unique identifier for each individual per frame.
  • Bounding Box: Pixel coordinates (xmin, ymin, xmax, ymax) for each annotated person.
  • Clip Type: Indicates real-world or cinematic origin of video clips.
Video naming follows the “[source identifier]_[sequence number]” format, where the identifier is the YouTube ID or film name, and the sequence number indicates its temporal position. This structured metadata approach ensures efficient integration between frame-level and video-level annotations, effectively supporting downstream vision-language research in microgravity environments. For comprehensive documentation of the dataset metadata, please refer to the resources provided on our GitHub and Hugging Face repositories.
Appendix F Supplementary for Human Action Recognition (HAR) Task
F.1 Dataset Partitioning
The dataset was partitioned at the video level following two primary criteria: (1) ensuring coverage of all action classes through greedy selection, and (2) proportional random splitting of remaining videos into training, validation, and test subsets (70:10:20 ratio). This approach avoids label leakage by maintaining video-level annotation consistency within subsets. Table 6 provides a summary of the sample-level and video-level distributions across splits.
Table 6: Train/val/test split statistics for the HAR task in MicroG-4M. Each cell reports both the absolute count and its corresponding percentage (%).
Split Samples (# / %) Video Clips (# / %)
Train 9,266 (69.93%) 3,331 (69.99%)
Val 1,329 (10.03%) 475 (9.98%)
Test 2,656 (20.04%) 953 (20.03%)
Total 13,251 (100.00%) 4,759 (100.00%)
F.2 Dataset Statistics
Table 7 summarizes the distribution of three broad action types, namely Object Manipulation, Person Interaction, and Person Movement, across the entire dataset, cinematic subset, and real video subset. Object Manipulation constitutes the largest category, particularly within real videos (39.42%), followed by Person Interaction and Person Movement categories.
Table 8 presents statistics regarding the number of persons annotated per video. Single-person videos dominate the dataset, particularly in real footage (92.41%), while multi-person annotations are comparatively more frequent in cinematic sources.
Additionally, Table 9 provides a detailed breakdown of 50 fine-grained action classes. The most frequent actions across both subsets include ‘stand’, ‘carry/hold object’, and ‘talk to self/person’, whereas actions such as ‘climb’, ‘take photo’, and ‘kiss person’ are comparatively rare. Notably, the distribution of action labels exhibits a pronounced long-tail pattern, consistent with Zipf’s law commonly observed in naturally occurring datasets, indicating the realistic and representative nature of the collected action annotations.
Table 7: Label type distribution across the full dataset, cinematic subset, and real video subset. Each cell shows the number of labels followed by its proportion (%).
Label Type All Videos Movies Real Videos
Object Manipulation 4,986 (37.60%) 1,416 (38.78%) 3,788 (39.42%)
Person Interaction 4,288 (32.34%) 1,198 (32.81%) 2,950 (30.70%)
Person Movement 3,987 (30.07%) 1,037 (28.40%) 2,872 (29.89%)
Table 8: Distribution of the number of persons per video in MicroG-4M. Each cell shows the number of videos followed by its proportion (%).
Persons per Video All Videos Movies Real Videos
Single Person (1) 3,983 (83.69%) 816 (61.26%) 3,167 (92.41%)
Two Persons (2) 623 (13.09%) 395 (29.65%) 228 (6.65%)
Three or More ($\geq$3) 153 (3.22%) 121 (9.09%) 28 (0.94%)
Table 9: Complete summary of action class distribution across all videos, movies, and real videos (sorted by action ID).
ID Action Name Count
All Videos Movies Real Videos
1 bend/bow (at waist) 26 14 12
3 crouch/kneel 20 2 18
5 fall down 10 1 9
6 get up 25 4 21
7 jump/leap 20 0 20
8 lie/sleep 17 4 13
9 martial art 18 0 18
10 run/jog 12 0 12
11 sit 252 239 13
12 stand 3218 698 2520
14 walk 369 75 294
17 carry/hold object 3126 549 2577
20 climb (e.g., mountain) 1 1 0
22 close (door/box) 13 4 9
24 cut 9 0 9
26 dress/undress clothing 31 9 22
27 drink 43 5 38
28 operate spaceship 20 16 4
29 eat 45 2 43
30 enter 68 5 63
34 hit object 33 3 30
36 lift/pick up 188 10 178
38 open (window/door) 32 13 19
41 play musical instrument 3 0 3
43 point to object 323 2 321
45 pull object 32 19 13
46 push object 24 8 16
47 put down 138 7 131
48 read 15 14 1
56 take photo 2 0 2
57 text/look at cellphone 7 0 7
58 throw 4 0 4
59 touch object 353 136 217
60 turn screwdriver 17 9 8
61 watch TV/unspecified 346 316 30
62 work on computer 110 67 43
63 write 3 3 0
64 fight/hit person 27 26 1
65 give/serve object 46 20 26
66 grab person 41 31 10
67 hand clap 3 3 0
68 hand shake 4 2 2
69 hand wave 140 44 96
70 hug person 16 13 3
72 kiss person 1 1 0
74 listen to person 148 135 13
76 push person 3 2 1
78 take object from person 15 10 5
79 talk to self/person $31$ 504 2627
80 watch person 713 625 88
F.3 Per-Class AP Results
Figures 7 and 8 display the per-class average precision (AP) performance of models fine-tuned on MicroG-4M and AVA datasets, respectively, and evaluated on the MicroG-4M test set. Fine-tuning on MicroG-4M generally achieves higher per-class AP scores, highlighting improved model adaptation to microgravity-specific action patterns. In contrast, AVA fine-tuned models exhibit significant performance degradation, particularly on fine-grained or microgravity-specific actions such as ‘operate spaceship’, ‘float’, or posture-related actions (‘bend/bow’). These differences underscore the critical need for specialized fine-tuning datasets such as MicroG-4M to address the unique challenges posed by microgravity environments in action recognition tasks.
Figure 7: Per-Class AP after Fine-Tuning on MicroG-4M, evaluated on the MicroG-4M test set. All models shown in this figure correspond to those listed in Table 1 of the main text.
Figure 8: Per-Class AP after Fine-Tuning on AVA, evaluated on the MicroG-4M test set. All models shown in this figure correspond to those listed in Table 2 of the main text.
F.4 Cross-Dataset VQA Density and Design Rationale
To contextualize our choice of six QA pairs per 3-second clip (2 QA/s), we compare MicroG-4M with representative video-VQA corpora in terms of clip granularity, QA density, and question types. This unified view allows controlled discussion of annotation budget and reasoning load across datasets with heterogeneous clip lengths.
Table 10: Cross-dataset VQA density and scope. Statistics are compiled from the original dataset reports. “QA/sec” denotes (Avg. QA/Clip)/(Clip length in seconds). This comparison contextualizes our choice of six QA per 3-second clip in MicroG-4M.
Dataset Clips QA pairs Avg. QA/Clip Clip Len. (s) QA/sec QA Types
MSVD-QA Xu et al. (2017) 1,970 50,505 25.64 10 2.56 Wh-type
MSRVTT-QA Xu et al. (2017) 10,000 243,680 24.37 15 1.62 Wh-type
TGIF-QA Jang et al. (2017) 56,720 103,919 1.83 3 0.61 Task-based
TVQA Lei et al. (2018) 21,793 152,545 7.00 76 0.09 Wh-type + Temporal
ActivityNet-QA Yu et al. (2019) 5,800 58,000 10.00 180 0.06 Motion, Spatial, Temporal
MovieQA Tapaswi et al. (2016) 6,771 6,462 0.95 200 0.004 Story comprehension
VideoQA Yang et al. (2003) 18,100 174,775 9.66 45 0.21 Wh-type + Yes/No
MicroG-4M (Ours) 1,238 7,428 6.00 3 2.00 Wh-type, Foreground/Background, Fine-/Coarse-motion, Identity, Temporal, Causal
MicroG-4M targets short clips at high QA density (2.00 QA/s), comparable to MSVD-QA and exceeding other widely used corpora. A fixed per-clip budget supports semantic diversity while avoiding redundancy typical of long-form settings. Our QA taxonomy balances foreground actions, spatial context, entity/attribute grounding, and temporal/causal reasoning, and explicitly includes an unanswerable option to reduce hallucination.
Because question generation protocols and scoring schemes vary across datasets, cross-corpus density should be interpreted as a comparability aid rather than a difficulty metric. The choice of six QA pairs per 3-second segment is thus motivated by controllability: it preserves clip-level coverage and evaluation stability while enabling apples-to-apples studies of domain shift in microgravity, independent of confounds from variable clip durations or QA volumes.